He sample of argon has a volume of 735 ml as a pressure of 1.20 ATM and a temperature of 112 degrees C what is the volume of the gas in millimeters on a pressure and temperature of the gas sample are changed to the following if the amount of gas Remains the Same?
A. 658 mm HG and 281 k
B. 0.55 ATM and 75 degrees C
C. 15.4 ATM and -15 degrees C
1
Expert's answer
2017-11-13T09:07:06-0500

(P_1∙V_1)/T_1 =(P_2∙V_2)/T_2

Before solving, convert T to kelvin, always in gas laws. If needs be, convert mm to atm by dividing mmHg / 760mmHg/atm. Then solving equations
A: P1=1.2 atm, V1=735 ml, T1=273+112 K, T2=281 K, P2=658 mmHg= 0.86579 atm so
V2=(P1*V1*T2)/(P2*T1) = (1.2*735*281)/( 0,87*385)= 679.8 ml
B: P1=1.2 atm, V1=735 ml, T1=273+112 K, T2=273+75 K, P2=0.55 atm, so
V2=(P1*V1*T2)/(P2*T1) = (1.2*735*348)/(0.55*385)= 1449.5 ml
C: P1=1.2 atm, V1=735 ml, T1=273+112 K, T2=273+(-15) K, P2=15.4 atm, so
V2=(P1*V1*T2)/(P2*T1) = (1.2*735* 258.15)/(15.4*385)= 38.4 ml

Answer: A=679.8 ml, B=1449.5 ml, C=38.4 ml